solve

7/x = 13/55

Answer by jim_thompson5910(35256)

7/x = 13/55

7*55 = 13x

385 = 13x

13x = 385

x = 385/13

Answer: x = 385/13

--------------------------------------------------------------------------------------------------------------

If you need more help, you can email me at jim_thompson5910@hotmail.com

To find other ways to contact me, you can also visit my website: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

--------------------------------------------------------------------------------------------------------------